# Edge densities of drawings of graphs with one forbidden cell

### v2

@@ Abstract
A connected topological drawing of a graph divides the plane into a number of cells. The type of a cell $c$ is the cyclic sequence of crossings and vertices along the boundary walk of $c$. For example, all triangular cells with three incident crossings and no incident vertex share the same cell type. When a non-homotopic drawing of an $n$-vertex multigraph $G$ does not contain any such cells, Ackerman and Tardos [JCTA 2007] proved that $G$ has at most 8n208n-20 edges, while Kaufmann, Klemz, Knorr, Reddy, Schröder, and Ueckerdt [GD 2024] showed that this bound is tight.
In this paper, we initiate the in-depth study of non-homotopic drawings that do not contain one fixed cell type $\mathfrak{c}$, and investigate the edge density of the corresponding multigraphs, i.e., the maximum possible number of edges. We consider non-homotopic as well as simple drawings, multigraphs as well as simple graphs, and every possible type of cell. For every combination of drawing style, graph type, and cell type, we give upper and lower bounds on the corresponding edge density. With the exception of the cell type with four incident crossings and no incident vertex, we show for every cell type $\mathfrak{c}$ that the edge density of $n$-vertex (multi)graphs with $\mathfrak{c}$-free drawings is either quadratic in $n$ or linear in $n$. In most cases, our bounds are tight up to an additive constant.
@@ 1 Introduction
Many fundamental classes of graphs are defined by admitting a drawing without a forbidden configuration or pattern. This includes planar graphs (forbidding two crossing edges), as well as many beyond-planar graph classes such as $k$-planar graphs (forbidding an edge with $k+1$ crossings), quasiplanar graphs (forbidding three pairwise crossing edges), fan-planar graphs (forbidding an edge crossed by two independent edges [^1] The actual definition is more complex but irrelevant here.) and many others. Planar and beyond-planar graphs are of eminent importance, both from a practical and a theoretical point of view. Nevertheless, many fundamental properties are still not sufficiently well understood. One of the most fundamental questions in this area concerns the edge density of a beyond-planar graph class $\mathcal{G}$, that is, the maximum possible number of edges of any $n$-vertex graph in $\mathcal{G}$. There is a long history of edge density results, ranging from $k$-planar graphs [pachGraphsDrawnFew1997, pachImprovingCrossingLemma2006, ackermanTopologicalGraphsMost2019], $k$-quasiplanar graphs [ackerman_quasiplanar, ackermanMaximumNumberEdges2009], $k$-bend RAC graphs [didimoDrawingGraphsRight2011, angeliniRACDrawingsGraphs2020], $(k,l)$-grid free graphs [pachTopologicalGraphsNo2005], and more to the recently introduced Density Formula [kaufmann_density_formula].
@@
Simple Non-homotopic             Construction˜88n288n-28       [ackerman_quasiplanar]8n208n-20
- Simple            Construction˜67n307n-30          Theorem˜6.17n207n-20
[Uncaptioned image]-cell Multi Non-homotopic             Construction˜59n189n-18 -
Simple Non-homotopic           Construction˜46n126n-12 $\binom{n}{2}$
- Simple           Construction˜46n126n-12
@@
[Uncaptioned image]-cell Multi Non-homotopic           Construction˜46n126n-12           Theorem˜4.16n126n-12
- Simple           Construction˜46n126n-12           Theorem˜4.16n126n-12
@@
Table 1: Results on edge densities of graphs admitting a drawing without one fixed type of cell. Upper bounds hold for all $n\geq 4$, while lower bounds are proven for infinitely many values of $n$.
@@
We start in Section˜2 with relevant definitions and some statements that we will use throughout the paper. Then, in Section˜3, we show that most cell types can be avoided in simple drawings of $K_{n}$ and that all cell types can be avoided in arbitrary drawings of $K_{n}$. Constituting the main body of this work, we address the remaining cell types in Sections 4 to 7, with quasiplanarity discussed jointly with [Uncaptioned image]-cells in Section˜6. Finally, in Section˜8, we consider the graph classes $\mathcal{G}_{\mathfrak{c}}$ induced by forbidding cell types $\mathfrak{c}$ whose boundary does not contain any crossings in more detail. We conclude this work with some open problems in Section˜9.
@@ 2 Preliminaries: Drawings, Cells, and Cell Types
Throughout this paper, we consider finite graphs with no loops but possibly parallel edges. To make the (potential) presence of multiedges more explicit, we sometimes use the term multigraph. Similarly, when a graph has no parallel edges, it is called a simple graph. A drawing $\Gamma$ of a multigraph $G=(V,E)$ maps vertices to pairwise distinct points and edges to curves connecting the corresponding vertices. For convenience, we mostly consider drawings on the sphere $\mathbb{S}^{2}$ to avoid a special treatment of the unbounded cell. As customary, we require that no edge passes through a vertex, any two edges have only finitely many points in common, each being a common endpoint or a proper crossing, and that no three edges cross in the same point. A drawing $\Gamma$ is connected if the image of the map $\Gamma$ is connected in $\mathbb{S}^{2}$.
@@
A drawing is simple if any two edges have at most one point in common, which is either a common endpoint or a crossing. As this rules out parallel edges, only simple graphs admit simple drawings. A lens is a region of $\mathbb{S}^{2}$ bounded by exactly two parts of edges. This could be two parallel edges, two edges crossing twice, or two crossing adjacent edges. We call a lens empty if its interior contains neither a crossing nor a vertex. Then a drawing is non-homotopic if it does not contain any empty lens.
@@
Consider any fixed connected drawing $\Gamma$ of some multigraph $G$. The crossings split each edge into a number of edge segments. An outer edge segment is incident to a vertex, while an inner edge segment starts and ends with a crossing. The cells of $\Gamma$ are the connected components of $\mathbb{S}^{2}$ after the removal of all vertices and edges. For a cell $c$, we denote by $v(c)$ the number of vertices, by $e(c)$ the number of edge segments, and by $e_{\rm in}(c)$ the number of inner edge segments on its boundary, counted with multiplicity. The size of $c$ is the sum of the former two: $\left\lVert c\right\rVert\coloneqq e(c)+v(c)$. We denote by $\mathcal{C}$ the set of all cells, and set $\mathcal{C}_{k}=\{c\in\mathcal{C}\colon\left\lVert c\right\rVert=k\}$ and $\mathcal{C}_{\geq k}=\{c\in\mathcal{C}\colon\left\lVert c\right\rVert\geq k\}$.
@@
Since $\Gamma$ is assumed to be a connected drawing of $G$, the boundary of every cell is connected. The type of a cell $c$ is the cyclic sequence of incidences with edge segments, crossings, and vertices along the boundary of $c$. We emphasize that a cell type $\mathfrak{c}$ only determines the form of incidence (edge segment, vertex, or crossing) and not the participating edge, vertex, or crossing of $G$. In particular, different incidences might be with the same edge segment, crossing, or vertex. We use pictograms with the size of the cell inscribed (such as [Uncaptioned image] or [Uncaptioned image]) to concisely denote cell types of small size. For convenience, let a cell type $\mathfrak{c}$ also denote the number of all $\mathfrak{c}$-cells in a given drawing. If no cell in $\Gamma$ has type $\mathfrak{c}$, then $\Gamma$ is called $\mathfrak{c}$-free.
@@
Let us conclude the preliminaries with an observation from [kaufmann_density_formula] and two lemmas, which we will need in the course of this paper.
@@
[Kaufmann, Klemz, Knorr, Reddy, Schröder, and Ueckerdt [kaufmann_density_formula]] Let $\Gamma$ be any non-homotopic connected drawing of some multigraph $G$ with at least three vertices. Then,
@@
The next two lemmata will help prove upper bounds for [Uncaptioned image]-free drawings in Section˜4 and [Uncaptioned image]-free drawings in Section˜6. Lemma˜2.1 below relates the number of vertices in a graph to the sizes of cells in a drawing of the graph. The statement already appears implicitly in various edge density proofs via discharging (e.g., [ackermanMaximumNumberEdges2009, arikushiGraphsThatAdmit2012, ackermanTopologicalGraphsMost2019]) and as a key ingredient in the proof for the Density Formula [kaufmann_density_formula, Lemma 3.6]. We thus state it without proof [^3] As one of many ways to prove it, for example, subtracting the Density Formula with $t=0$ from the Density Formula with $t=1$ gives the result..
@@
The next lemma gives a bound on the number of cells of certain types in terms of the number of inner edge segments in large cells of non-homotopic drawings. Similar statements have been implicitly used in prior work such as [ackerman_quasiplanar, ackermanTopologicalGraphsMost2019, kaufmann_density_formula].
@@ 3 Drawings of Complete Graphs
Figure 1: A simple drawing of $K_{n}$ (here $n=8$) containing only cells of size at most 55, and one cell of size $2n$.
@@
Next, we show that [Uncaptioned image]-cells can also be avoided in simple drawings of $K_{n}$. Note that any drawing without uncrossed edges will do, see for example [harborthEdgesWithoutCrossings1974]. For completeness, we present our own construction.
@@
Let $n\geq 8$ be even and start with a straight-line drawing of $K_{n}$ with vertices in convex position. Then redraw the edges of length $3$ as depicted in Figure˜2, so as to cross the edges of length $1$ (on the convex hull). Further redraw every other edge of length $2$ to be uncrossed in the outer cell. Finally, redraw all edges of length $n/2$ to be pairwise crossing in the outer cell. For $n$ odd, take the construction for n1n-1 even, add the last vertex in the central cell, and connect it to all other vertices with straight-line segments.
@@
For any cell type $\mathfrak{c}$ and infinitely many values of $n$, the complete graph $K_{n}$ admits a $\mathfrak{c}$-free drawing.
@@
4 Drawings without [Uncaptioned image]-Cells
In this section, we consider the [Uncaptioned image]-cell. In particular, we show that non-homotopic [Uncaptioned image]-free drawings have only linear edge density. Note that by Construction˜3, the restriction to non-homotopic drawings is essential.
@@ 4 Drawings without [Uncaptioned image]-Cells
Fix a [Uncaptioned image]-free drawing of an $n$-vertex multigraph with $n\geq 3$. First, let us compute for each cell $c\in\mathcal{C}$ the quantity 3e(c)+2v(c)123e(c)+2v(c)-12, and in case $c\in\mathcal{C}_{\geq 5}$ compare it with $e_{\rm in}(c)$.
@@
Theorem˜4.1 gives an upper bound of 6n126n-12 on the edge density of [Uncaptioned image]-free non-homotopic multigraphs. As the next construction shows, this bound is tight, even for simple drawings.
@@
Figure 4: Left: A simple Refer to caption-free and Refer to caption-free drawing with 6n126n-12 edges. Right: A non-homotopic Refer to caption-free drawing of a multigraph with 9n189n-18 edges.
@@
5 Drawings without [Uncaptioned image]-Cells
In this section, we briefly discuss the [Uncaptioned image]-cell. From Construction˜4, we immediately obtain a 6n126n-12 lower bound for the edge density of simple graphs that admit a [Uncaptioned image]-free drawing. A slight variation of this construction yields a better bound for non-homotopic drawings.
@@ 5 Drawings without [Uncaptioned image]-Cells
Thus, the obtained drawing is [Uncaptioned image]-free. Non-homotopicity is apparent since only the pairs of parallel edges form lenses, which are non-empty. Finally, the drawing contains 9n129n-12 edges, since for each of the 3n63n-6 edges of $G_{0}$, two edges are added.
By Construction˜5, the edge density of [Uncaptioned image]-free non-homotopic multigraphs is at least 9n189n-18. For simple [Uncaptioned image]-free graphs, it is at least 6n126n-12 by Construction˜4. However, we have no non-trivial upper bound; see also Problem˜9.1 in Section˜9.

### v3

@@ Abstract
A connected topological drawing of a graph divides the plane into a number of cells. The type of a cell $c$ is the cyclic sequence of crossings and vertices along the boundary walk of $c$. For example, all triangular cells with three incident crossings and no incident vertex share the same cell type. When a non-homotopic drawing of an $n$-vertex multigraph $G$ does not contain any such triangular cell, Ackerman and Tardos [JCTA 2007] proved that $G$ has at most 8n208n-20 edges, while Kaufmann, Klemz, Knorr, Reddy, Schröder, and Ueckerdt [GD 2024] showed that this bound is tight.
In this paper, we initiate the in-depth study of $\mathfrak{c}$-free drawings, that is, drawings that do not contain any cell of one fixed cell type $\mathfrak{c}$, and investigate the edge density of the corresponding graphs, i.e., the maximum possible number of edges. We consider non-homotopic as well as simple drawings, multigraphs as well as simple graphs, and every possible cell type $\mathfrak{c}$. For every combination of drawing style, graph type, and cell type, we give upper and lower bounds on the corresponding edge density. With the exception of the cell type with four incident crossings and no incident vertex, we show for every cell type $\mathfrak{c}$ that the edge density of $n$-vertex (multi)graphs with $\mathfrak{c}$-free drawings is either linear in $n$ or superlinear in $n$. In most cases, our bounds are tight up to an additive constant. We further consider the question which simple graphs admit a simple drawing without some given cell type(s). For the class of cell types that are not incident to any crossing, we give a complete characterization of all simple graphs that admit a simple drawing without any such cell.
@@ 1 Introduction
Many fundamental classes of graphs are defined by admitting a drawing without a forbidden configuration or pattern. This includes planar graphs (forbidding two crossing edges), as well as many beyond-planar graph classes such as $k$-planar graphs (forbidding an edge with $k+1$ crossings), quasiplanar graphs (forbidding three pairwise crossing edges), fan-planar graphs (forbidding an edge crossed by two independent edges [^1] The actual definition is more complex but irrelevant here.) and many others. Planar and beyond-planar graphs are of eminent importance, both from a practical and a theoretical point of view [didimoSurveyBeyondGraphDrawing2020]. Nevertheless, many fundamental properties are still not sufficiently well understood. One of the most fundamental questions in this area concerns the edge density of a beyond-planar graph class $\mathcal{G}$, that is, the maximum possible number of edges of any $n$-vertex graph in $\mathcal{G}$. There is a long history of edge density results, ranging from $k$-planar graphs [pachGraphsDrawnFew1997, pachImprovingCrossingLemma2006, ackermanTopologicalGraphsMost2019], $k$-quasiplanar graphs [ackerman_quasiplanar, ackermanMaximumNumberEdges2009], $k$-bend RAC graphs [didimoDrawingGraphsRight2011, angeliniRACDrawingsGraphs2020], $(k,l)$-grid free graphs [pachTopologicalGraphsNo2005], and more to the recently introduced Density Formula [kaufmann_density_formula].
@@
[…]
- Simple           Cons. 8 based on [ackerman_quasiplanar]7n307n-30       [ackerman_quasiplanar]7n207n-20
@@
[…]
Table 1: Results on edge densities of $n$-vertex graphs admitting a drawing without one fixed type of cell. Upper bounds hold for all $n\geq 4$, while lower bounds are proven for infinitely many values of $n$. Bounds written in bold are shown in this work.
@@
We start in Section˜2 with relevant definitions and some statements that we will use throughout the paper. Then, in Section˜3, we show that most cell types can be avoided in simple drawings of $K_{n}$ and that all cell types can be avoided in arbitrary drawings of $K_{n}$. Constituting the main body of this work, we address the remaining cell types in Sections 4 to 6, with quasiplanarity discussed jointly with [Uncaptioned image]-cells in Section˜5. Finally, in Section˜7, we consider the graph classes $\mathcal{G}_{\mathfrak{c}}$ induced by forbidding cell types $\mathfrak{c}$ whose boundary does not contain any crossings in more detail. We conclude this work with some open problems in Section˜8.
@@ 2 Preliminaries: Drawings, Cells, and Cell Types
Throughout this paper, we consider finite graphs with no loops but possibly parallel edges. To make the (potential) presence of multiedges more explicit, we sometimes use the term multigraph. Similarly, when a graph has no parallel edges, it is called a simple graph. A drawing $\Gamma$ of a (potentially disconnected) multigraph $G=(V,E)$ maps vertices to pairwise distinct points and edges to curves connecting the corresponding vertices. For convenience, we mostly consider drawings on the sphere $\mathbb{S}^{2}$ to avoid a special treatment of the unbounded cell. As customary, we require that no edge passes through a vertex, any two edges have only finitely many points in common, each being a common endpoint or a proper crossing, and that no three edges cross in the same point. A drawing $\Gamma$ is connected if the image of the map $\Gamma$ is connected in $\mathbb{S}^{2}$.
@@
A drawing is simple if any two edges have at most one point in common, which is either a common endpoint or a crossing. As this rules out parallel edges, only simple graphs admit simple drawings. A lens in a drawing is a region of $\mathbb{S}^{2}$ bounded by exactly two parts of edges. These could be two parallel edges, two edges crossing twice, or two crossing adjacent edges. We call a lens empty if its interior contains neither a crossing nor a vertex. A drawing is non-homotopic if it does not contain any empty lens.
@@
Consider any fixed connected drawing $\Gamma$ of some multigraph $G$. The crossings split each edge into a number of edge segments. An outer edge segment is incident to a vertex, while an inner edge segment starts and ends with a crossing. The cells of $\Gamma$ are the connected components of $\mathbb{S}^{2}$ after the removal of all vertices and edges. For a cell $c$, we denote by $v(c)$ the number of vertices, by $e(c)$ the number of edge segments, and by $e_{\rm in}(c)$ the number of inner edge segments on its boundary. Here, vertices and edge segments may be counted multiple times; as often as they appear when walking along the boundary of $c$. The size of $c$ is the sum of the former two quantities: $\left\lVert c\right\rVert\coloneqq e(c)+v(c)$. We denote by $\mathcal{C}$ the set of all cells, and set $\mathcal{C}_{k}=\{c\in\mathcal{C}\colon\left\lVert c\right\rVert=k\}$ and $\mathcal{C}_{\geq k}=\{c\in\mathcal{C}\colon\left\lVert c\right\rVert\geq k\}$.
@@
Since $\Gamma$ is assumed to be a connected drawing of $G$, the boundary of every cell is connected. The type of a cell $c$ is the cyclic sequence of incidences with edge segments, crossings, and vertices along the boundary of $c$. We emphasize that a cell type $\mathfrak{c}$ only determines the form of incidence (edge segment, vertex, or crossing) and not the participating edge, vertex, or crossing of $G$. In particular, different incidences might be with the same edge segment, crossing, or vertex. We use pictograms with the size of the cell inscribed (such as [Uncaptioned image] or [Uncaptioned image]) to concisely denote cell types of small size. For convenience, we let these pictograms also denote the number of cells of that type in a given drawing. If no cell in $\Gamma$ has type $\mathfrak{c}$, then $\Gamma$ is called $\mathfrak{c}$-free.
@@
Let us conclude the preliminaries with an observation by Kaufmann, Klemz, Knorr, Reddy, Schröder, and Ueckerdt [kaufmann_density_formula], and two lemmas, which we will need in the course of this paper.
@@
[[kaufmann_density_formula, Observation 2.2]] Let $\Gamma$ be any non-homotopic connected drawing of some multigraph $G$ with at least three vertices. Then
@@
The next two lemmata will help prove upper bounds for [Uncaptioned image]-free drawings in Section˜4 and [Uncaptioned image]-free drawings in Section˜5. Lemma˜2.1 below relates the number of vertices in a graph to the sizes of cells in a drawing of the graph. The statement already appears implicitly in various edge density proofs via discharging (e.g., [ackermanMaximumNumberEdges2009, arikushiGraphsThatAdmit2012, ackermanTopologicalGraphsMost2019]) and as a key ingredient in the proof for the Density Formula [kaufmann_density_formula, Lemma 3.1]. We thus state it without proof [^3] As one of many ways to prove it, for example, subtracting the Density Formula with $t=0$ from the Density Formula with $t=1$ gives the result..
@@
The next lemma gives a bound on the number of [Uncaptioned image]-cells and [Uncaptioned image]-cells in terms of the number of inner edge segments in large cells of non-homotopic drawings. Similar statements have been implicitly proven in prior work [ackerman_quasiplanar, ackermanTopologicalGraphsMost2019].
@@
According to [kaufmann_density_formula, Lemma 3.6], for the set $S_{\rm in}$ of all inner edge segments we have
@@
On the other hand, every inner edge segment appears twice on the boundary of some cell. As the only cells of size at most four are [Uncaptioned image]-, [Uncaptioned image]-, and [Uncaptioned image]-cells (cf. Section˜2), it follows
@@ 3 Drawings of Complete Graphs
Figure 1: A simple drawing of $K_{n}$ (here $n=8$) containing only cells of size at most 55, and one cell of size $2n$. The numbers in the figure indicate the sizes of the different kinds of cells.
@@
Let $n\geq 8$ be even and start with a straight-line drawing of $K_{n}$ with vertices as equidistant points on a circle. By the length of an edge we refer to the distance of its endpoints in circular order along the circle. Redraw the edges of length $3$ as depicted in blue in Figure˜2, so as to cross the edges of length $1$ (on the convex hull). Note that this change retains simplicity. Further, redraw every other edge of length $2$ to be uncrossed in the outer cell (red edges in Figure˜2). Finally, redraw all edges of length $n/2$ to be pairwise crossing in the outer cell (green edges in Figure˜2). This again retains simplicity. For $n$ odd, take the construction for $n^{\prime}=n-1$, add the last vertex in the central cell, and connect it to all other vertices with straight-line segments. The resulting drawing $\Gamma$ is simple. Further, since every edge is crossed at least once, $\Gamma$ is also [Uncaptioned image]-free .
@@
For any cell type $\mathfrak{c}$ and any $n\geq 8$, the complete graph $K_{n}$ admits a $\mathfrak{c}$-free drawing.
@@
Locally replace each crossing with the pattern in Figure˜3. In the resulting drawing, each cell incident to a crossing is part of such a pattern and therefore of size $2$ or at least $6$.
@@
4 Drawings without [Uncaptioned image] or [Uncaptioned image]-Cells
In this section, we consider the [Uncaptioned image]- and [Uncaptioned image]-cell. We start by showing that non-homotopic [Uncaptioned image]-free drawings have only linear edge density. Note that by Construction˜3, the restriction to non-homotopic drawings is essential. This result can also be derived from the proof of a statement on straight-line drawings with crossing angle restrictions [dujmovicLargeAngleCrossing2011, Theorem 4.2].
@@ 4 Drawings without [Uncaptioned image] or [Uncaptioned image]-Cells
Fix a [Uncaptioned image]-free drawing of a connected $n$-vertex multigraph with $n\geq 3$. First, let us compute for each cell $c\in\mathcal{C}$ the quantity 3e(c)+2v(c)123e(c)+2v(c)-12, and in case $c\in\mathcal{C}_{\geq 5}$ compare it with $e_{\rm in}(c)$.
@@
Theorem˜4.1 gives an upper bound of 6n126n-12 on the edge density of [Uncaptioned image]-free non-homotopic drawings of multigraphs. As the next construction shows, this bound is tight, even for simple drawings. This construction also provides our best lower bound for the edge density of simple [Uncaptioned image]-free drawings.
@@
Figure 4: Left: A simple Refer to caption-free and Refer to caption-free drawing with 6n126n-12 edges. Right: A non-homotopic Refer to caption-free drawing of a multigraph with 9n189n-18 edges. On both sides, $n=12$.
@@
A slight variation of this construction yields a better bound for non-homotopic [Uncaptioned image]-free drawings.
@@
Thus, the obtained drawing is [Uncaptioned image]-free. Non-homotopicity is apparent since only the pairs of parallel edges form lenses, which are non-empty. Finally, the drawing contains 9n129n-12 edges, since for each of the 3n63n-6 edges of $G_{0}$, two further edges are added.
By Construction˜5, the edge density of [Uncaptioned image]-free non-homotopic multigraphs is at least 9n189n-18. For simple [Uncaptioned image]-free graphs, the edge density is at least 6n126n-12 by Construction˜4. However, we have no non-trivial upper bound; see also Problem˜8.1 in Section˜8.